OpenStudy (anonymous):

the ramp below is used to move crates of fruit to loading docks of different heights. when the horizontal distance AB is 12 meters, the height of the loading dock, BC, is 4 meters. what is the height of the loading dock? 12 meters 8 meters 9 meters 15 meters

OpenStudy (anonymous):

this is the ramp

Vocaloid (vocaloid):

hint: set up a proportion height of small triangle/base of small triangle = height of big triangle/base of big triangle

OpenStudy (anonymous):

@ganeshie8 @Nnesha

OpenStudy (anonymous):

thats the part i need help on

Vocaloid (vocaloid):

4/12 = x/(12+15) solve for x

OpenStudy (anonymous):

would it be 8 meters? @Vocaloid

Vocaloid (vocaloid):

nope, check your math again

Vocaloid (vocaloid):

start by cross-multiplying

OpenStudy (anonymous):

12?

Vocaloid (vocaloid):

please stop guessing and solve for x

OpenStudy (anonymous):

im not guessing

Vocaloid (vocaloid):

then show me your math

Vocaloid (vocaloid):



OpenStudy (anonymous):

4 x 12 = 44 12 x 15 = 180 exactly how r u supposed

OpenStudy (anonymous):

48 sorry

OpenStudy (anonymous):

27 x 4 is 108 so 12 time what is x makes no sense

OpenStudy (anonymous):



OpenStudy (anonymous):

i swear it has to be 8

OpenStudy (anonymous):

@Allenismyname @tapha @TheSmartOne @alekos

OpenStudy (tapha):

srry im here

OpenStudy (anonymous):

so basically 4(27/4)

OpenStudy (anonymous):

4 12 = x 12+15 Step 1: Cross-multiply. 4 12 = x (12+15) 4*(12+15)=x*(12) 108=12x Step 2: Flip the equation. 12x=108 Step 3: Divide both sides by 12. 12x 12 = 108 12 x=9

OpenStudy (anonymous):

i mean 12 V.v

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

oh i multiplied that part thats why i got 8 ok thanks

OpenStudy (anonymous):

i agree ^

OpenStudy (anonymous):

Sure, thing!

OpenStudy (tapha):

8